OpenStudy (abmon98):

a curve has equation 12/3-2x i) find dy/dx a point moves along this curve. As the point passes through A, the x-coordinate is increasing at a rate of 0.15 units per second and the y-coordinate is increasing at a rate of 0.4 units per second find the possible coordinates of A

OpenStudy (abmon98):

dy/dx=-12/(3-2x)^2

OpenStudy (amistre64):

that is a complex function correct?

OpenStudy (abmon98):

That's a rate of Change Differentiation

OpenStudy (abmon98):

question

OpenStudy (amistre64):

12/3-2x i ^ it has an imaginary component right?

OpenStudy (amistre64):

or thats just i) to indicate part 1 of a set of questions

OpenStudy (abmon98):

dy/dx=8/3

OpenStudy (amistre64):

y = 12/3-2x y (3-2x)= 12 3y-2xy= 12 3y' - 2x'y - 2xy'= 0 3(.4) - 2(.15)y - 2(.4)x= 0 1.2 - .3y - .8x= 0 12 = 3y + 8x

OpenStudy (abmon98):

i guess my derivative is incorrect

OpenStudy (amistre64):

yours was: dy/dx=-12/(3-2x)^2 you forgot a chain rule, multiply by -2

OpenStudy (abmon98):

12(3-2x)^-1=-2(-12)/(3-2x)^2

OpenStudy (amistre64):

thats better yes, so 40/15 = 24/(3-2x)^2

OpenStudy (abmon98):

i solve x=3 or 0

OpenStudy (amistre64):

yes

OpenStudy (abmon98):

thanks :)

OpenStudy (amistre64):

yw